OpenStudy (josedavid):

i need help please in 1 question

OpenStudy (josedavid):

Find the perimeter of the shape below: A four sided figure ABCD is shown. A is located at negative 4, 3. B is located at 0, 2. C is located at 2, 4. D is located at 0, negative 3 17.6 21.4 26.7 32.9

OpenStudy (josedavid):

@nuttyliaczar

OpenStudy (josedavid):

@nathan917 @nikeboi101 @NetflixAndNo @pooja195

OpenStudy (josedavid):

@agent0smith @Agl202 @bia_gonzalex

OpenStudy (automaticloveletter):

you switch from bio to history to math real quick son

OpenStudy (josedavid):

this is math

OpenStudy (agent0smith):

Use the distance formula to find the length of each side. Then add up all the sides.

OpenStudy (radar):

Review this link: http://www.purplemath.com/modules/distform.htm Then use the distance formula for each segment. After each distance is determined, add them (all four) and that will be the perimeter.

OpenStudy (josedavid):

is d

OpenStudy (josedavid):

guys

OpenStudy (radar):

Quit guessing and start working the problem.

OpenStudy (josedavid):

is d

OpenStudy (radar):

Check your work and see if there is an error.

OpenStudy (radar):

What did you get for the length of the segment AB ?

OpenStudy (radar):

Help is what you will get. But it looks like you just want an answer rather than learn how to solve.

OpenStudy (radar):

Just in case you return I will show you how to get the distance of the segment AB. The endpoints for segment AB are (-4,3) and (0,2). Using the distance formula which was explained in the link I furnished, and also in your study material we would have this equation:\[D = \sqrt{(-4-0)^{2 }+(3-2)^{2}}\]\[\sqrt{(-4)^{2}+1^{2}}=\sqrt{17}\]

OpenStudy (radar):

Theres one, now just three more to go. Do the rest in a similar manner, the perimeter will be their sum (all four segments.